How you can find an arithmetic sequence that is related to the sequence of perfect cubes, 1, 8, 27, 64,125,...?
Answer by KMST(5400)

Strange question.
The difference between consecutive terms of the sequence of perfect cubes are
7, 19, 37, 61, ....
The differences between consecutive differences are
19-7=12
37-19=18
61-37=24
and so on.
The differences of the differences, 12, 18, 24, ... form an arithmetic sequence.
Maybe that is the expected answer.